If the design of a solid shaft based on the maximum-distortion-energy leads to c (AMP® +37° where is the radius of the shaft, as soon as possible
Answer #1

Reactions at A and B in 6 RA+ RB = 420 +420 840 ebf. Taking moment about A; 420 x 1 + 420X 4 = RoxG Ro 350 lbf RA 840 - 350 a120 + 300 420lbf 300+120= 4201bf D 바이터 3 ft ya aft RB = 350 lbf 700 efft RA= 490 lbf - 490 lbfft Bending moment diagram B 700For unit conversion all = 77 ksi 778 6.895 MPa = 530.896 MPa 700 lbf.ft - 700x 1.36 Nm a g49.07 Nm Nm T= 180 lbf ft 180 X1.36